### v2

# Non-Linear Scoring Model for Translation Quality Evaluation

Abstract
Analytic Translation Quality Evaluation (TQE), based on Multidimensional Quality Metrics (MQM), has traditionally used a linear error-to-penalty scale calibrated to a reference sample size of 1000100020002000 words. However, linear extrapolation is known to bias judgment on samples of a different size than the reference, over-penalizing short samples and under-penalizing long ones, resulting in persistent misalignment with expert intuition.
Building on the previously introduced Multi-Range framework [GladkoffEtAl2024], this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length. This approach presents empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size.
This premise is supported by psychophysical and cognitive evidence, including the Weber–Fechner law and Cognitive Load Theory, which explain why the impact of additional errors diminishes perceptually but grows cognitively with scale. We propose a two‑parameter model
$E(x)=a\ln\!\bigl(1+bx\bigr),\quad a,b>0,$
that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step. The model yields an explicit interval (in closed form) within which the linear approximation stays within $\pm 20\%$ relative error, and it drops into existing evaluation workflows unchanged apart from a dynamic tolerance function.
It improves interpretability, fairness, and Inter-Rater Reliability (IRR) across both human and AI-generated translations. By reinforcing and operationalizing a perceptually valid scoring paradigm, this work advances the field toward a more accurate, scalable, and cognitively grounded translation quality evaluation. The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation congruent with human judgment. Implementation in CAT/LQA systems and implications for human and AI-generated text evaluation are discussed.
1 Introduction
Analytic Translation Quality Evaluation (TQE) is widely recognized as the gold standard method for assessing translation quality with high precision. It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014], with proper training of annotators. Once the annotations are in place, a scoring model is applied to produce a quantifiable quality assessment.
In practical terms, most industry scoring models operate under a linear assumption: the number of allowed errors scales proportionally with the word count of the evaluated sample. This model typically normalizes errors per 1000 or 2000 words to compare them against a client-defined threshold, e.g., no more than five minor errors per 1000 words. The normalized error density is then used to make a pass-or-fail decision or compute a quality score.
This method works reasonably well for evaluation sample sizes close to the reference size—that is, the standard word count used as a baseline for comparing results (typically 1,000–2,000 words). However, problems arise when evaluating smaller or larger samples.
Clients, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality. Waddington’s comparative studies [waddington1999, waddington2001] show that additive, points-off models alone do not align with holistic judgments; combining analytic and holistic perspectives improves reliability. This suggests that perceived translation quality is not a linear function. Evidence from other domains points in the same direction. In speech and video quality, international standards report top-end ’compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic]. Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ’showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450]. Large-scale MT evaluations similarly show that expert, context-aware severity judgments yield system rankings that diverge from simple averages, supporting non-additive scoring [freitag2021]. And in usability engineering, a single ’catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994]. Together, these applied literatures reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
In points‑off situations, the linear rule makes long samples appear better than experts judge them to be by raising the acceptance threshold in direct proportion to length.
A recent quote from a large institutional client illustrates this point:
’Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them. We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient. The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula. For example, we might feel that if a translation sample is about one page, a single major mistranslation error is already enough to judge it as failing; in a seven-page sample, however, seven such errors are far more than we would be willing to tolerate. Instead, we would prefer to fail the sample already at three or four major errors. This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.’
Another client confirmed these findings and chose to avoid the issue by only evaluating samples of exactly 1000 words. However, in real-world operations, such precisely consistent evaluation sample sizes are rare. Evaluation samples often vary significantly in length due to the nature of the content, deadlines, or operational constraints.
Despite its known practical shortcomings, the assumption of linear error scaling remains entrenched in industry practice—partly due to implicit conventions and the replication of established procedures, and partly because people often subconsciously apply linear extrapolation to numerical judgments, including those about error tolerance. Yet when evaluators (and therefore also content users) assess samples without relying on mental calculation, their judgments reliably deviate from linear expectations.
This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length. The present study argues that as the evaluated text gets longer, the maximum number of allowable errors grows, but more and more slowly—equivalently, the perceptionally allowed errors per 1,000 words decline as word count increases. We capture this pattern with a logarithmic dependency.
Earlier work [GladkoffEtAl2024] introduced a logarithmic scorecard. This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory); we then provide a simple, easily calibrated model for use in current evaluation tools.
2 Background
2.1 Analytic TQE and MQM
Analytic Translation Quality Evaluation (TQE) refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications. Unlike holistic methods, which rely on a general impression of quality, analytic TQE requires systematic annotation of specific error types, each classified under predefined categories.
The most prominent and widely adopted framework for such structured annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014]. MQM provides a hierarchical error typology that is flexible and extensible, with high-level error types like Accuracy and Linguistic Conventions (formerly called Fluency) and more granular error subtypes such as Mistranslation, Omission, Grammar, and Spelling. Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the client’s preference.
Once the annotation phase is complete, a scoring model is applied to an evaluation sample (a defined word count from the job), converting the annotated errors into a numeric score or a Pass/Fail verdict. The MQM framework is currently the most detailed and accurate approach to analytic translation-quality evaluation in professional practice. Every MQM-based metric has two inseparable parts: (i) an error typology that specifies which error types are annotated (with a severity level such as minor/major/critical), and (ii) a scoring model that maps those annotations to severity-weighted penalty points and a quality score. For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words.
Refinements to either component therefore directly advance the state of the art in translation-quality measurement.
2.2 Linear Scoring Models and Their Limits
The dominant approach to scoring in translation quality evaluation is the linear model. Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size (typically 1,000–2,000 words) so that assessments are comparable across jobs of different lengths. For example, a limit of five minor errors one point each per 1,000 words implies 2.5 penalty points in a 500-word sample and 10 in a 2,000-word sample (and if a major error carries a $5\times$ severity multiplier, a single major error consumes the full 5-point allowance for a sample of 1000 words).
Although this proportional scaling appears objective, it fails however to reflect how human content users actually perceive quality. Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450]; however, this tolerance wanes as the document length grows further and the errors begin to compound. For instance, as noted above, a single major error is often sufficient to fail a one-page sample; in a seven-page sample, however, evaluators would fail it at far fewer than seven major errors—even though a linear rule would imply proportional tolerance. Such intuitive resistance reveals that human perception of quality is fundamentally non-linear.
Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
  • micro-range (<250<250 words): For very short samples (under 250 words), traditional quality scoring methods are unreliable. As shown in [gladkoff2021], quality scores for such brief samples exhibit extremely wide confidence intervals—meaning the range of possible true scores is so large that the measurement is imprecise—making individual measurements meaningless for decision-making.
    This limitation is particularly important for automatic Quality Estimation (QE), especially with the practice of sentence- or segment-level scoring. Recent industry findings confirm that these fine-grained predictions do not align with how humans perceive quality. Bane[baneATA2025] highlighted the repeated failures of sentence- or segment-level QE in high-volume translation, where automated scores fail to identify all major, and even critical, errors, making the method unreliable. As discussed in [gladkoff2025substack], attempting to derive meaningful scores from very small samples or isolated sentences or segments simply does not work.
    Instead, for micro-range samples, Statistical Quality Control (SQC) methods offer a more statistically robust and risk-aware framework for interpreting quality signals, explicitly framing the producer’s and consumer’s risk associated with very limited observations.
  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a client-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another linear model is needed; a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
The present study concentrates on the entire span of statistically reliable measurements in both the meso-range and the macro-range and supplies both psychophysical and empirical evidence for adopting a logarithmic scoring model across these ranges. Here “statistically reliable” is meant in the technical sense—i.e., estimates with reasonably narrow confidence intervals —the usual sampling-uncertainty bands around an estimate that shrink as sample size grows. [^1] For intuition, a simple 95% Wald interval for a proportion is $p\pm 1.96\sqrt{p(1-p)/n}$ [gladkoff2021]. For small samples or extreme proportions, this estimate is known to undercover; a Wilson (score) interval—or its closely related Agresti–Coull “plus-four” approximation—provides markedly better coverage [AgrestiCoull1998]. Wilson 95% CI: $\frac{\hat{p}+\frac{z^{2}}{2n}\pm z\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}}{4n^{2}}}}{1+\frac{z^{2}}{n}},$ with z=1.96z\!=\!1.96; the Agresti–Coull “plus-four” uses $\tilde{p}=(x+\tfrac{z^{2}}{2})/(n+z^{2})$ and $\tilde{p}\pm z\sqrt{\tfrac{\tilde{p}(1-\tilde{p})}{n+z^{2}}}$ [AgrestiCoull1998].
When linear scoring is applied indiscriminately across all ranges, it produces systematic friction. Clients and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples. Some organizations respond by limiting evaluations to a fixed 1,0001{,}000-word sample, but this workaround is operationally restrictive and fails to address the underlying perceptual misalignment. Because the width of a confidence interval shrinks roughly like $1/\sqrt{n}$—with an additional finite-population correction $\sqrt{(N-n)/(N-1)}$ when the evaluated sample is a large share of the document—variation in the sampled word count can change those intervals considerably. [^2] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
The persistent use of linear models—despite long-recognized shortcomings—stems largely from the absence of simple, theoretically grounded alternatives. This paper aims to fill that gap by presenting a non-linear scoring model and providing empirical and theoretical support for it, grounded in psychophysics and cognitive-load research.
3 Empirical Evidence of Non-Linearity
To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large enterprise clients. Each of these organizations manages multilingual translation programs with substantial daily volumes and relies on mature, metric-based TQE systems.
3.1 Survey Design
The extended calibration questionnaire described in [GladkoffEtAl2024] was reused, asking survey respondents to state acceptable error counts for multiple evaluation sample sizes without performing linear extrapolation. This method again proved effective in eliciting a latent non-linear response curve.
A detailed calibration questionnaire was designed to capture Quality Managers’ intuition about acceptable error tolerance—operationalized as the maximum number of severity-weighted penalty points (APT) they would still pass—across a range of evaluation sample sizes.
Importantly, specific measures were taken to neutralize the influence of the standard linear model, which respondents are accustomed to using in their daily work. Without intervention, professionals often default to mental linear math—extrapolating tolerances from known thresholds. To counteract this bias, the clients’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
The questionnaire asked the following questions:
  • Intuitively, how many minor errors would you accept in a 1-page sample?
  • Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages?
  • For a 3-page text, how many minor errors seem tolerable—just based on your experience?
  • Don’t calculate—just estimate: how many minor errors would you allow in 4 pages?
  • Based on gut feeling, how many minor errors would still be okay in a 5-page sample?
  • Without applying your formal model, what number of minor errors seems acceptable in 10 pages?
  • Without referencing formulas, what minor error count would intuitively feel acceptable for 20 pages?
Elicitation controls.
To reduce linear anchoring, page sizes were presented in randomized order without showing prior answers; two page sizes were repeated at the end to probe intra‑rater stability. Respondents were instructed explicitly not to prorate from known thresholds and to answer from professional judgment alone.
The questionnaire deliberately focused on minor errors for two key reasons. First, higher-severity errors (major and critical) are too infrequent for reliable statistical analysis. Second, as many studies note, their impact on quality perception is significantly different and more pronounced. This specific focus is practical, as standard industry practice allows the severity of any error to be expressed as an equivalent number of minor errors.
While clients used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.
3.2 Results
The most striking finding was a clear non-linear pattern in error tolerance. Freed from the habit of linear extrapolation, respondents indicated that while they would accept more total errors in longer texts, their tolerance grew far more slowly than the sample size. Instead, the tolerance curve rose steeply at first and then gradually flattened. This pattern was consistent across all three clients, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types.
Figures 13 illustrate the raw responses. For each, the horizontal axis represents the sample size in pages, and the vertical axis shows the maximum number of minor errors that the respondent deemed acceptable. In one case, both minor and major error series were collected.
Figure 1: Client 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
Figure 2: Client 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
Figure 3: Client 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
None of the clients’ tolerance curves followed a straight line. Although each client had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve.
The respondents themselves were surprised by this result. When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ’acceptable’ was non-linear. They became aware that their tolerance per page was higher in shorter samples and lower in longer ones, contrary to what linear extrapolation would dictate.
This phenomenon aligns with everyday experience. In very short samples, a single error can be jarring and dominate the user’s entire impression of quality. In longer documents, however, the dynamic shifts: while the perceived impact of any one mistake diminishes, the overall experience is instead degraded by the steady accumulation of minor issues, which erodes trust and creates a sense of general sloppiness.
This degradation impacts more than just a pleasant "reading experience"; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly. Ultimately, these findings confirm that the linear scoring model does not reflect how humans judge translation quality: acceptability is not a linear function of error count, because the perceived impact of an error depends on how much text has already been processed—early mistakes loom larger, while later ones add less per unit but accumulate as the sample grows.
Consistent with Fig. 7 in [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span.
Based on this consistent pattern, this paper hypothesizes that human tolerance to translation errors grows logarithmically with the sample size. The rest of this paper explores the theoretical underpinnings of this phenomenon and proposes a formal scoring model that better aligns with real-world perceptions.
From Figures 1–3 it is already clear that the data points do not fall on a straight line. The logarithmic curve, in contrast, follows the shape of the expert judgments closely.
Quantitative comparison of linear vs. logarithmic fits (technical note)
For readers who want the formal evidence behind Fig. 3: we fit both a straight line and a logarithmic curve to the same elicited points and compared their goodness-of-fit using standard metrics. Smaller is better for SSE (sum of squared errors) and RMSE (root-mean-square error); larger is better for $R^{2}$ (coefficient of determination). AIC (Akaike Information Criterion) and BIC (Bayesian Information Criterion) are model-selection scores that penalize extra parameters; lower is better.
$\mathrm{SSE}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\qquad\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
where $n$ is the number of points and $k$ the number of fitted parameters. As Table 1 shows, the logarithmic model has dramatically smaller SSE/RMSE, much higher $R^{2}$, and much lower AIC/BIC, so its advantage is not an artifact of the extra parameter but a closer match to the elicited judgments. These all confirm, quantitatively, that the log model explains about $95\%$ of the variance ($R^{2}\!=\!0.94$), while the linear model explains very little ($R^{2}\!=\!0.04$), and that the log model’s advantage is not merely due to having one extra parameter.
In other words, the visual impression is also statistically decisive.
Below is the quantitative comparison of linear vs. logarithmic fits on the elicited points.
To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes. This is what the values in Table 1 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better). Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC. This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can. In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
Logarithmic $a\ln(1+bx)$ $3.353$ $0.590$ $1.551$ $0.471$ $0.945$ 6.550-6.550 6.658-6.658
Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7, k=2k{=}2 (log) and k=1k{=}1 (linear). Smaller is better.
Table 1: Goodness-of-fit statistics for the Client 1 minor-error dataset ($x$ in pages; $E$ in minor points). Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit. Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).
The standard indicators reported in Table 1 show how well a curve follows the observed data points. They illustrate why the logarithmic model offers a markedly superior fit than the linear one. [^3] SSE (sum of squared errors): the total squared distance between the curve and the data (smaller is better). RMSE: the average size of that distance (smaller means the model sits almost on top of the data). $R^{2}$: the fraction of variability explained by the curve ($R^{2}\approx 1$ means excellent, $R^{2}\approx 0$ means poor). AIC and BIC: compact comparison measures that reward better fit but penalize extra parameters (lower is better). Thus, even without going into formulas, the table makes visible that the non-linear model is overwhelmingly better at reflecting real human tolerance. When multiple elicited points are available, we recommend weighted least squares (App. B), down‑weighting sparse or less reliable page sizes. Using the delta method in App. B, we obtain approximate prediction standard errors and plot shaded confidence bands around the fitted curve.
4 Psychophysical and Cognitive Foundations
Empirical findings from these client interviews show that perceived translation quality is not a simple linear function of error count or text length. Instead, human judgment follows deep-seated perceptual and cognitive regularities documented since the nineteenth century [weber1834, fechner1860, boring1950]. This paper draws on two complementary frameworks—classical psychophysics and modern Cognitive Load Theory—to motivate a logarithmic scoring model.
4.1 Weber–Fechner Law and Logarithmic Perception
Psychophysics has long recognized that sensation grows non-linearly with stimulus intensity. In his 1834 monograph on weight discrimination, Weber reported that just-noticeable difference (JND) is a constant fraction of the baseline stimulus: “The smallest perceptible increase of weight bears a fixed ratio to the weight already lifted[weber1834, p. 152]. Three decades later, Fechner integrated Weber’s proportionality into a mathematical model of subjective sensation, arguing that “if equal stimulus ratios yield equal sensation increments, then sensation must be the logarithm of stimulus magnitude[fechner1860, vol. 1, p. 55]. This logarithmic relation—now known as the Weber–Fechner law—was the first quantitative statement of a compressive perceptual scale; an authoritative English translation is provided in [fechner1966].
Intuition.
Applied to translation quality evaluation, the law would predict that the subjective penalty of an additional error is highest when the text is nearly error-free and decreases as errors accumulate. In other words, the first mistranslation in a short paragraph feels much more jarring than the fifteenth typo in a 20-page report.
Subsequent psychophysical research refined, but did not overturn, this insight. Stevens showed that many modalities follow a power function $S=k\,I^{n}$ with exponents $n<1$ for brightness, loudness, and heaviness [stevens1957]. Although $n=0$ is not itself logarithmic, the power-law family $I^{n}$ is asymptotically log-like:
$\lim_{n\to 0}\frac{I^{\,n}-1}{n}\;=\;\ln I.$
Over finite dynamic ranges, exponents in the $0.2\text{–}0.4$ band produce curves that closely approximate a logarithmic profile. Over typical stimulus ranges, power exponents $n\in[0.2,0.4]$ are empirically indistinguishable from a log on plots like ours, so a logarithmic tolerance is the parsimonious choice here. Comprehensive reviews such as [gescheider1997] and the signal detection synthesis of [green1966] confirm that compressive (sublinear) mapping is the norm in vision, audition and touch.
Hence, long before today’s quality assessment challenges, experimental psychology had already established a robust theoretical framework: Human judgment converts objective intensities into logarithm-like internal scales. Treating translation errors as discrete ’stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
4.2 Cognitive Load and Accumulated Impact of Errors
Whereas psychophysics explains the decrease in sensitivity to additional stimuli, cognitive psychology accounts for how these stimuli collectively strain mental resources. The theory of limited working memory, formalized as Cognitive Load Theory (CLT), provides a parallel foundation for non-linear perception of translation quality.
CLT was introduced by Sweller in the late 1980s as a way to explain why instructional design must respect the cognitive limitations of learners [sweller1988]. It builds on earlier findings such as Miller’s famous observation that the number of discrete items a person can hold in short-term memory is limited to approximately seven, plus or minus two [miller1956]. As Sweller notes, “Our capacity to process novel information is extremely limited, and when this limit is exceeded, learning and comprehension suffer[sweller1988, p. 257].
In translation quality evaluation, the same principle applies: each identified error acts as a disruption that requires attentional and interpretive resources. One or two errors may be brushed aside, but a buildup of errors, however minor individually, begins to saturate the evaluator’s cognitive capacity. This leads to increased frustration, reduced comprehension, and a degraded overall impression of quality.
Importantly, this degradation does not occur linearly. Just as in perceptual scaling, there is a threshold effect: the cumulative weight of repeated small errors eventually causes a sharp drop in processing fluency. Studies in instructional design and reading comprehension (e.g. [mcw2020]) have documented how disfluencies compound non-linearly and trigger avoidance behaviors or reduced engagement.
Cognitive Load Theory also provides a plausible explanation for the empirical finding that users tend to expect stricter quality in longer documents, even when the overall error density remains stable [waddington1999, waddington2001]. This reflects a protective mechanism: anticipating that accumulated disruptions will erode comprehension, readers subconsciously raise the standard for what they consider acceptable in large-scale content.
Together, these cognitive effects mirror and reinforce the non-linear sensitivity modeled by the Weber–Fechner framework. Although psychophysics focuses on the perceptual scaling of individual stimuli, Cognitive Load Theory models the impact of cumulative disruption. Both lead to the same conclusion: the relationship between perceived translation quality and the raw error count (that is, the objective number of penalty points) is fundamentally non-linear.
4.3 Perception and Cognition Converge on Logarithmic Scoring
Together, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically. A calibrated logarithmic curve therefore offers the most parsimonious way to reconcile (i) diminishing marginal impact of individual errors with (ii) the accelerating cognitive cost of cumulative errors. This synthesis underpins the non-linear scoring model proposed in the next section.
5 Proposed Non-Linear Scoring Model
This paper adopts the terminology introduced in [GladkoffEtAl2024], distinguishing between three classes of scoring models.
  • Raw Score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC}$, with $\alpha$ a unit factor (e.g., $\alpha=1000$ for “per 1,000 words”). This uses the sample’s penalty rate and is not a calibrated/anchored score.
  • Calibrated Linear: This scoring model maps a client’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs.
  • Calibrated Non-Linear: A model that replaces the straight-line rule with a logarithmic tolerance curve. Relative to a linear rule anchored at the reference size, the model is more lenient for short samples and stricter for long ones; total allowance still grows with length, but at a slowing rate.
Given both the empirical evidence and the theoretical justifications presented, this paper proposes a more nuanced application of scoring models rather than a complete replacement of the linear approach. While the traditional linear model is often considered sufficient for the meso-range (texts up to 3,000 words), our findings indicate this is precisely where its limitations can be most acute. For example, as shown later in this paper, a linear model calibrated for a 1,000-word reference sample loses significant accuracy (over 20%) when applied to a 2,000-word sample.
Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in §5.4). Both the traditional raw model and the calibrated linear model remain viable tools, but their use should be restricted to scenarios where evaluation sample sizes consistently hover very close to the specific word count for which the models were calibrated. 444‘Calibration’ for the raw score simply means choosing the client’s pass/fail cut-point on the raw penalty-rate scale. Let $r_{\text{thr}}$ be the maximum allowed error rate (e.g., 5 points per 1,000 words). A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes if and only if $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$. No calibration mapping to a 0–100 quality scale is involved. Note that some teams report a 0–100 “raw score” by first scaling the penalty rate to a per-1,000-word unit and then computing $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this rescaling still inherits the limitations of raw scoring. To obtain comparable scores across evaluation sample sizes, a calibrated model (linear near the reference size or non-linear for wider ranges) is required.
This section introduces the mathematical form of the proposed logarithmic model, explains how it can be calibrated, and demonstrates its practical advantages.
By introducing a Non-Linear Calibrated scoring model, this paper extends the MQM scoring apparatus itself, enabling it to represent perceptual tolerance that linear models cannot capture. Because the same logarithmic function adapts seamlessly across evaluation sample sizes, language pairs, and content domains, an MQM-based Non-Linear Calibrated metric becomes congruent with holistic (document-level) expert judgment—turning MQM into a practically universal translation quality evaluation system. [^5] Other scoring models may be devised for specific applications. Importantly, however, an exponential ladder of severity weights is not the same thing as a non-linear scoring model in the sense used here. Severity weights (even exponential ones) change how individual errors are combined within a fixed sample; they do not by themselves make the pass/fail tolerance as a function of sample size non-linear. For example, the ATA certification rubric is an analytic, points-off scheme tailored to short, fixed-length exam translations. Its severity scale is steep (many exponential levels), so higher-severity errors count far more than minor ones, but aggregation is still essentially additive and the exam length is fixed; there is no tolerance curve $E(x)$ that varies with $x$. Similarly, SAE J2450 (automotive) assigns fixed category weights, producing a non-uniform penalty distribution but still summing points linearly within a sample [saeJ2450]. (One could reproduce similar effects by choosing MQM type weights accordingly.) Thus ATA and SAE J2450 introduce severity-level non-uniformities, not the sample-size non-linearity addressed in this paper. They are effective for their narrow contexts, but they do not provide the general, non-linear tolerance function $E(x)$ needed to keep scores comparable when evaluation size varies widely. A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope]. LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution. In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size. The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
5.1 Limitations of Linear Extrapolation
Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size. For example, if a client accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc. However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
As noted above, many institutional clients describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound. In practice, however, evaluators do not extend that allowance proportionally. When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors. In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length. What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole. This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
5.2 Mathematical Model
This paper proposes the following logarithmic function to model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$, expressed in words:
$E(x)\;=\;a\,\ln(1+b\,x),\qquad a>0,\;b>0.$ (1)
Throughout, $\ln$ denotes the natural logarithm.
If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged. This unit invariance ensures figures shown in pages and formulas written in words are consistent. The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens. This curve shape effectively combines both theoretical expectations and observed data.
To clarify on the form of this function, the intercept-free form $E(x)=a\,\ln(1+bx)$ is deliberately used, which enforces $E(0)=0$ and allows curvature control via $b$, not the Excel spreadsheet “logarithmic trendline” $c+k\ln x$. The latter introduces a free intercept and does not pass through the origin, so it cannot be anchored to calibration points in the same way.
5.3 Anchoring and Calibration Procedures
This paper distinguishes between two steps in adapting a scoring model to operational use:
  • Anchoring defines the maximum acceptable error limit for a standard amount of text. This tolerance is expressed in penalty points based on a client’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
  • Calibration takes the absolute penalty point total and maps it onto a normalized, more easily interpretable 0-100 quality scale for use in dashboards or QA reports.
Feasibility (two‑point). Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
$r:=\frac{E_{1}}{E_{0}},\qquad\rho:=\frac{x_{1}}{x_{0}},\qquad g(b):=\frac{\ln(1+bx_{1})}{\ln(1+bx_{0})}.$
For $b>0$, $g$ is continuous and strictly monotone, with $\lim_{b\downarrow 0}g(b)=\rho$ and $\lim_{b\uparrow\infty}g(b)=1$. Hence a unique $b^{\star}>0$ exists if and only if
$\boxed{\ \min\{1,\rho\}\;<\;r\;<\;\max\{1,\rho\}\ }.$
Intuitively, the longer sample must permit more errors but sublinearly; the shorter must permit fewer errors but not in direct proportion to length.
The model can be easily calibrated using two known tolerance points. Suppose we have:
  • The primary anchor point: This is the main tolerance defined during the Anchoring step, such as a maximum of 5 errors ($E_{0}=5$) for a reference sample size of 1,000 words ($x_{0}=1000$).
  • A second, empirical data point: This point is derived from observing the client’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
Two‑point calibration (numerical).
Let $r=E_{1}/E_{0}$ and define $f(b)=\ln(1+bx_{1})-r\,\ln(1+bx_{0})$. Under the feasibility condition above, $f$ has exactly one root $b^{\star}>0$ by the intermediate value theorem. Near $b=0$, $f(b)\approx b(x_{1}-rx_{0})$, so the sign at the origin is $\mathrm{sgn}(x_{1}-rx_{0})$. As $b\to\infty$, $f(b)\sim(1-r)\ln b+\ln x_{1}-r\ln x_{0}$; if $r>1$ the limit is -$\infty$, and if $r<1$ it is $\infty$, guaranteeing a sign change and thus a unique root. Solve $f(b)=0$ with a robust one‑dimensional method (e.g., bisection/Brent); then set
$a^{\star}=\frac{E_{0}}{\ln(1+b^{\star}x_{0})}.$
See  Appendix A. Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
Illustrative values.
Feasibility.
Ensure $r=E_{1}/E_{0}$ lies strictly between $1$ and $\rho=x_{1}/x_{0}$, i.e., $\min\{1,\rho\}<r<\max\{1,\rho\}$, before solving for $b$.
Solving $f(b)=0$ yields $b^{\star}\approx 2.880\times 10^{-3}$ and $a^{\star}\approx 3.688$, giving the calibrated curve
$E(x)\;=\;3.688\,\ln\!\bigl(1+0.00288\,x\bigr).$
Fallback for multiple points. If more than two tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ are available—or if the two chosen points violate the feasibility condition—estimate $(a,b)$ by constrained least squares:
$\min_{a>0,\;b>0}\ \sum_{i=1}^{n}\Bigl[E_{i}-a\,\ln\!\bigl(1+b\,x_{i}\bigr)\Bigr]^{2}.$
Given any $b_{0}>0$, a fast 1‑D search over $b$ with the closed‑form $a(b)=\dfrac{\sum_{i}E_{i}\,\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}$ works well (then refine both jointly if desired).
This approach allows the model to be tailored to specific content types, organizational standards, or language pairs while still operating under a common mathematical framework. For instance, the error tolerance can be precisely defined for each case, ranging from extremely strict for software UI text, to moderately strict for marketing copy and documentation, and significantly more lenient for internal communications or user-generated content.
For small $b$, $\ln(1+b\,x)\approx b\,x$, so the model reduces locally to a linear tolerance rule; this explains why a calibrated linear slope performs acceptably only in the vicinity of its reference sample size.
If $x$ is expressed in pages with $W$ words per page, the same curve is obtained by replacing $b$ with $W\cdot b$; $a$ is unchanged.
Figure 4: Three interpolations. The spreadsheet curve $c+k\ln x$ is shown for comparison only; calibration uses $E(x)=a\ln(1+bx)$, which passes through the origin and is anchorable to tolerance points.
5.4 Near-Reference Fidelity Interval for Linear Approximation (±20%)
In practice, a linear rule anchored at a reference size $x_{\mathrm{ref}}$ is often used for convenience:
$E_{\mathrm{lin}}(x)\;=\;E_{\log}(x_{\mathrm{ref}})\,\frac{x}{x_{\mathrm{ref}}},\qquad E_{\log}(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr).$
To quantify where this linear approximation is acceptable, we define the ±20% fidelity interval around $x_{\mathrm{ref}}$ as the set of $x>0$ such that
$\left|\frac{E_{\mathrm{lin}}(x)}{E_{\log}(x)}-1\right|\leq 0.20.$
Writing $\varepsilon=0.20$ and $\alpha=\dfrac{\ln(1+b\,x_{\mathrm{ref}})}{b\,x_{\mathrm{ref}}}$, the two boundary points admit a closed form via the Derivation (sketch). Set $E_{\rm lin}(x)=(x/x_{\rm ref})\,E_{\rm log}(x_{\rm ref})$ and require $E_{\rm lin}(x)=(1\pm\varepsilon)E_{\rm log}(x)$. Writing $\alpha=\ln(1+bx_{\rm ref})/(bx_{\rm ref})$ and $y=1+bx$, we obtain $\alpha\ln y=(1\pm\varepsilon)(y-1)/y$. Rearranging gives $y\,e^{-\frac{1\pm\varepsilon}{\alpha}\,y}=e^{-\frac{1\pm\varepsilon}{\alpha}}$, so
$y=-\frac{1\pm\varepsilon}{\alpha}\,W_{-1}\!\Bigl(-\frac{\alpha}{1\pm\varepsilon}\,e^{-\alpha/(1\pm\varepsilon)}\Bigr),$
which yields the closed form in the next equation. The branch $W_{-1}$ selects the solution near $x_{\rm ref}$.
Lambert $W$ function:
$x_{\pm}\;=\;\frac{-\,\dfrac{1\pm\varepsilon}{\alpha}\,W_{-1}\!\!\left(\,-\frac{\alpha}{1\pm\varepsilon}\,e^{-\alpha/(1\pm\varepsilon)}\right)-1}{\,b\,}$,
where $x_{-}$ (with the “-” sign) gives the lower boundary (linear is $0.8\times$ the log curve) and $x_{+}$ (with the “++” sign) gives the upper boundary (linear is $1.2\times$ the log curve). [^6] The branch $W_{-1}$ is selected to obtain the solution near $x_{\mathrm{ref}}$.
Let’s look at the numerical intervals for the example calibration. Using the illustrative curve from §5.3 (a=3.688,b=0.00288a=3.688,\,b=0.00288), the ±20% fidelity intervals are:
  • $\boldsymbol{x_{\mathrm{ref}}=1000}$ words: $x\in[579,\;1460]$ words, width $\approx 881$ words (about $42\%$ to $46\%$ relative to $x_{\mathrm{ref}}$).
  • $\boldsymbol{x_{\mathrm{ref}}=2000}$ words: $x\in[1307,\;2747]$ words, width $\approx 1440$ words (about $34.7\%$ to $37.4\%$ relative to $x_{\mathrm{ref}}$).
These calculations illustrate how far one can move away from the reference sample size while still retaining acceptable measurement accuracy. The window is relatively narrow: if the reference is 1,000 words, then evaluating a 2,000-word sample already falls outside the ±20% fidelity band and will not be accurate; likewise, if the reference is 2,000 words, applying the same linear rule to a 1,000-word sample is equally unreliable.
In absolute terms, the ±20% window widens as the reference size increases (from about 0.88 k to about 1.44 k words), but in relative terms it becomes proportionally narrower. This quantifies the earlier point: a calibrated linear rule is only reliable near its anchoring point. Further away, it becomes too harsh for short samples and too lenient for long ones. When jobs regularly fall outside these fidelity intervals, the logarithmic model should be used directly, or the linear rule should be re-anchored to a more suitable reference size.
5.5 Visual comparison across three sample-size ranges
Figure 5: Linear model (blue) vs. human perception (PFT, red). Yellow bands indicate the $\pm 20\%$ fidelity zone around the linear anchor; outside these bands the linear rule over‑ or under‑estimates true tolerance.
Figure 5 visualizes this critical weakness of the linear model: Once the slope is fixed using a single calibration point, the linear model severely diverges from the actual human judgment at both lower and higher evaluation sample sizes. In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation wordcount).
The calibrated logarithmic curve is more permissive for short samples and stricter for long ones relative to an anchored linear rule; it increases with length but flattens, avoiding the linear model’s over‑leniency at large sizes.
This not only matches human content users’ expectations, but also reduces misclassification of borderline quality outputs.
Figure 6: Applicability across evaluation sample sizes. SQC (micro, $\lesssim\!250$ words), calibrated linear MQM (meso; reliable only within $\sim\!\pm 20\%$ of anchor), and non-linear MQM (macro). Vertical mark at $\sim\!250$ words flags the micro‑range where deterministic curves are unreliable. Y‑axis: error tolerance (penalty points).
Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics. In the micro-range (fewer than 250 words), evaluation outcomes are highly sensitive to single errors and statistical confidence is very low; thus, risk-based Statistical Quality Control (SQC) is most appropriate.
For the meso- and macro-ranges, the choice of an appropriate scoring model depends not on the absolute length of any single text, but on the expected range of fluctuation in evaluation sample sizes. This paper’s findings point to a crucial rule of thumb: if evaluation sample sizes are expected to vary by more than approximately 20% from the anchor point, the non-linear model is required for accuracy. A calibrated linear model is only a viable tool when evaluation sample sizes remain consistently within this tight  20% band around the anchor. Counter-intuitively, because the logarithmic curve flattens at large values, a linear model can be more applicable in the macro-range than the meso-range, provided its anchor point is also in that higher range.
Ultimately, however, for any operational scenario involving a wide variety of evaluated text lengths, the non-linear logarithmic model is the only option that ensures consistent and reliable accuracy. This is because a single linear extrapolation becomes unreliable as it moves further from its anchor point, whereas a logarithmic model remains accurate across both the meso- and macro-ranges.
Scenario Recommended model
EWC varies within $\pm 20\%$ around the anchor Calibrated linear model (re-anchor if the range shifts)
EWC routinely outside the $\pm 20\%$ band Use the non-linear logarithmic model directly
Very short samples ($\lesssim 250$ words) Statistical Quality Control (SQC); deterministic curves unreliable
Table 2: Operational guidance: which scoring model to use in practice.
Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
6 Practical Implications
To the extent that error tolerance is not linear — as suggested by both empirical findings and practitioner reports — the adoption of a logarithmic scoring model can reshape translation quality evaluation practices, quality management processes, and software tooling across the industry. Replacing the linearity assumption allows quality assessment to become more accurate, consistent, and aligned with human perception. Looking ahead, such a model might also serve as a reference framework for emerging AI-based document-level evaluations. Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (client-specific PT/MSV mapping).
6.1 Operational Guidelines for Quality Managers
One immediate benefit of the non-linear model is its ability to inform and standardize quality expectations across variable evaluation sample sizes. Quality managers can use the calibrated log function to define clear error tolerance thresholds for different content types or use cases, reducing ambiguity when evaluating translation outputs that differ in word count.
For example, a quality manager might specify: “Our reference wordcount is 2000 words with a tolerance of 7 minor errors. For evaluation sample sizes between 1500 and 3000 words, the standard linear approximation is acceptable. Outside that range, the logarithmic model must be used to adjust expectations accordingly.”
This shift allows quality managers and stakeholders to move beyond one-size-fits-all scoring rules and instead adopt quality scores that reflect actual cognitive effort and perceived severity. The result is more fair and interpretable evaluation outcomes, especially when comparing translation outputs across different content types or word counts.
7 Building your non-linear scorecard
Notation. The designations and symbols in this section follow the variable naming used in the ASTM working draft WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail.
7.1 Model and least-squares fit
We model the pass/fail tolerance as a function of sample size $x$ (words) by
$E(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr),\qquad a>0,\;b>0.$ (2)
Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$, estimate $(a,b)$ by least squares
$\min_{a>0,b>0}\;\sum_{i=1}^{n}\bigl[E_{i}-a\,\ln(1+b\,x_{i})\bigr]^{2}.$
Profiling out $a$ yields, for any fixed $b>0$,
$a(b)\;=\;\frac{\sum_{i}E_{i}\,\ln(1+b\,x_{i})}{\sum_{i}\ln^{2}(1+b\,x_{i})},\qquad\text{then minimize }S(b)=\sum_{i}\!\bigl[E_{i}-a(b)\ln(1+b\,x_{i})\bigr]^{2}\text{ in }b>0,$ (3)
[…]
If $x$ was fitted in pages with $W$ words/page and you now compute in words, only $b$ rescales: $b_{\text{words}}=b_{\text{pages}}/W$; $a$ is unchanged.
7.2 From tolerance to a calibrated score
For an evaluation of size $x=\text{EWC}$ (words), compute the allowed penalty points at the passing threshold
$E_{\text{allowed}}(x)\;=\;\hat{a}\,\ln\!\bigl(1+\hat{b}\,x\bigr).$ (4)
Let $\mathrm{APT}$ be the absolute penalty total from the MQM table. Define the quality fraction
$\mathrm{QF}(x):=1-\frac{\mathrm{APT}}{E_{\text{allowed}}(x)}.$ (5)
Map $\mathrm{QF}$ to the standard calibrated scale using Maximum Score Value (MSV), Passing Threshold (PT), and Defined Passing Interval $\mathrm{DPI}=\mathrm{MSV}-\mathrm{PT}$:
$\mathrm{OS}(x):=\mathrm{PT}+\mathrm{DPI}\,\mathrm{QF}(x),\qquad\mathrm{OS}_{\text{disp}}:=\min\{\mathrm{MSV},\,\max\{0,\mathrm{OS}(x)\}\}.$ (6)
The rating is Pass if $\mathrm{APT}\leq E_{\text{allowed}}(x)$, else Fail.
Monotonicity and clipping. Since $\mathrm{OS}(x)=\mathrm{PT}+\mathrm{DPI}\bigl(1-\mathrm{APT}/E_{\mathrm{allowed}}(x)\bigr)$ is affine in $\mathrm{APT}$, it decreases monotonically as $\mathrm{APT}$ grows; displaying $\mathrm{OS}_{\rm disp}=\min\{\mathrm{MSV},\max[0,\mathrm{OS}(x)]\}$ does not affect Pass/Fail, which always uses the un‑clipped inequality $\mathrm{APT}\leq E_{\mathrm{allowed}}(x)$.
7.3 Scorecard mapping
We now make the non-linear tolerance usable in everyday QA. The table below shows how to compute the scorecard fields from the evaluation word count (EWC), the fitted coefficients $\hat{a},\hat{b}$, and the severity-weighted penalty total (APT). The MQM error table itself is unchanged; the only change is that the allowed penalty $E_{\text{allowed}}(x)$ depends on the sample size.
Name Excel-style computation
Evaluation size $x$ (EWC) input
Model coefficients $\hat{a},\hat{b}$ calibrated from tolerance points (see §5.3)
Allowed penalty at size $x$, $E_{\mathrm{allowed}}$ = a*LN(1 + b*EWC)
Quality fraction QFQF $1-\frac{\mathrm{APT}}{E_{\mathrm{allowed}}}$
Calibrated score $\mathrm{OS}$ = PT + (MSV-PT) * QF
Displayed score $\mathrm{OS}_{\mathrm{disp}}$ = MIN(MSV, MAX(0, OS))
Decision margin DMDM $E_{\mathrm{allowed}}-\mathrm{APT}$
Pass/Fail = IF(APT <= Eallowed,"PASS","FAIL")
Table 3: Scorecard mapping: variables and Excel-style computations.
This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
Figure 7: Example of a real-world non-linear calibrated scoring model for documentation content type (solid yellow curve) plotted against the linear model with 10 allowed penalty points per 1000 words (dashed line). The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual client tolerance function; the linear line is drawn for comparison only.
Same-curvature retargeting (optional). If you want a new anchor $E(x_{\text{ref}})=E_{0}$ while maintaining curvature, keep $b$ and solve
$a\;=\;\frac{E_{0}}{\ln(1+b\,x_{\text{ref}})}.$
This re-anchors the tolerance without changing the curve shape.
Micro-range caution. Below $\sim$250 words, deterministic tolerance curves are statistically unreliable; in this sample size range methods of Statistical Quality Control must be used rather than score curves. Quantitatively, if the underlying minor‑error rate is $p$ per word and we observe only $n$ words, the expected count is $np$ with standard deviation $\sqrt{np(1-p)}\approx\sqrt{np}$, so the relative uncertainty scales as $1/\sqrt{np}$ and becomes huge when $np\approx 1$ (e.g., 1 expected error in 200 words at 5/1000).
Worked example (documentation content). Calibrated tolerance: $E(x)=3.688\,\ln(1+0.00288\,x)$. For $x=3000$ words, $E_{\mathrm{allowed}}(x)\approx 8.357$ points. If $\mathrm{APT}=7$, then $QF=1-7/8.357\approx 0.162$. With $\mathrm{PT}=80$, $\mathrm{MSV}=100$ ($\mathrm{DPI}=20$), $\mathrm{OS}\approx 83.25$ and $DM\approx 1.36$ (Pass). If $\mathrm{APT}=9$, then $QF\approx-0.077$, $\mathrm{OS}\approx 78.46$ and $DM\approx-0.64$ (Fail).
7.4 Implications for CAT Tools and LQA Automation
The change in scoring logic has clear consequences for software systems that implement translation quality evaluation, particularly CAT tools with integrated LQA modules or standalone dedicated LQA platforms.
In the traditional linear model, these tools calculate scores based on a fixed error density relative to a single reference word count. However, with a logarithmic tolerance function, the score calculation must dynamically adjust based on the actual sample size. This means LQA modules and dashboards need to:
  • Replace static pass/fail thresholds based on a single anchor point with dynamic tolerance functions, $E(x)$, that adjust based on the actual word count ($x$).
  • Visualize quality scores against a curved baseline that represents the tolerance function, not a simple flat line.
  • Provide evaluators with interactive feedback showing how close a sample’s error count is to its unique, size-adjusted pass/fail threshold.
This functionality is particularly valuable for machine translation, post-editing, and AI-generated content, as evaluation samples from these sources often vary widely in word count.
7.5 Improvement of a Unified Calibrated Quality Scale
One major challenge in TQE is ensuring that quality scores remain human-readable, interpretable, and comparable across different types of content, language pairs, and evaluation sample lengths. The linear calibrated model is a significant improvement over using raw scores with very small passing intervals, but even the linear calibrated model can distort comparability when applied to large or small samples.
With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, client-specific quality scale. Since the logarithmic tolerance accounts for sample-size effects, and calibration maps results to a common 0–100 (PT/MSV) scale, all evaluations land on the same company-wide quality scale—so scores from jobs of different lengths, language pairs and content types are directly comparable.
For example, replicating the English–German marketing case from Table 3 in [GladkoffEtAl2024], a 5,000-word translation with 23 minor errors passed the linear model, but was rejected by human reviewers. The logarithmic model set a threshold of 16 errors—correctly matching the expert decision.
This normalization also improves the validity of quality measurement results analytics across a portfolio of projects, enabling better tracking of supplier performance, trend analysis, and root-cause investigations.
7.6 Applications Beyond Human Translation
While the immediate application of this model is human translation quality evaluation, the underlying principles are equally relevant to assessing AI-generated content. Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation.
The model based on non-linear error tolerance provides a fairer and more realistic baseline for evaluating content. It avoids the common pitfalls of unduly penalizing shorter segments with one mistake or overlooking the accumulated minor issues in long, highly fluent AI-generated texts.
As organizations adopt more automated content generation, the ability to evaluate quality in a human-aligned, perceptually realistic manner becomes essential. The proposed scoring model provides a foundation for doing so.
7.7 Document-Level Evaluation and Future Automation
Translation quality never exists in a vacuum: every segment lives within the cohesive fabric of a document. A translation sample is therefore not a roll of toilet paper to be inspected centimeter by centimeter; its quality must be judged in the context of the larger text. The larger the sample relative to the document, the more faithfully that judgment reflects holistic properties such as cohesion, coherence, register, and terminological consistency.
Small extracts can still provide a rapid signal, but offer only a limited view of document quality. They cannot capture discourse-level phenomena, and—because error density interacts non-linearly with length—linear metrics break down outside their narrow reference window.
The MQM-based Non-Linear Calibrated model solves this by remaining simple and linear in the meso-range yet extending smoothly into the macro-range. As long as the evaluation sample is sufficiently large relative to the document, as recommended by the Multi-Range Framework[GladkoffEtAl2024]—it yields scores that are statistically reliable and consistent with the judgment of experts at the document level.
Of course, analytic TQE becomes increasingly expensive at large evaluation sample sizes. No current AI system can yet perform a complete analytic evaluation, but future models may acquire the ability to conduct formal TQE (full analytic MQM evaluation) over long contexts (document-scale samples), reliably capturing local errors (segment-level) and holistic properties (cohesion, coherence, terminology/style consistency, discourse). The logarithmic model proposed here is expressly designed to serve as the computational backbone for such future automation, ensuring that automated scoring remains perceptually valid across the full spectrum of document lengths.
7.8 Summary
In practical terms, the transition to a non-linear scoring model means:
  • Greater alignment between future automated document-level scoring systems and expert human judgment,
  • Improved fairness and consistency across evaluation samples of different lengths,
  • More intelligent quality evaluation systems producing evaluations closer to human perception,
  • Stronger foundations for evaluating both human and AI-generated translation.
This model equips both translation professionals and tool developers with a framework that respects the cognitive and perceptual realities of language evaluation, enhancing decision-making at every level of the translation quality pipeline.
The proposed non-linear model captures how translation quality is actually perceived by humans-especially when evaluation sample sizes vary. It is simple to implement, easy to calibrate, and theoretically grounded. By adopting non-linear scoring, evaluation systems can better align numeric scores with human judgment, increasing the reliability and fairness of translation quality evaluations, as well as increasing the reliability and fairness of quality evaluations of translations created by humans or AI-generated texts.
8 Limitations and Future Work
The following limitations and research directions qualify the benefits summarized in §7.8 and the conclusion.
While the findings and proposed non-linear quality evaluation model presented in this paper are supported by both empirical observation and theoretical grounding, there are important limitations that must be acknowledged. Addressing these will be essential for refining the model and validating its broader applicability.
8.1 Limited Survey Group Size
We conducted structured interviews with quality managers at three very large enterprise clients, and the results are corroborated by their internal practices. Despite the scale and maturity of these organizations’ multilingual translation programs, the interview sample is still small. Additional evidence was drawn from our own practice, where we exclusively use a non-linear calibrated scoring model due to its universality and accuracy. However, the tolerance data we used was gathered through expert elicitation and professional experience, rather than under blind experimental conditions—introducing subjectivity and potential sources of bias, including response bias.
Further research should involve a broader range of organizations—including LSPs, in-house localization teams, and public-sector translation departments—across different domains and language pairs. Ideally, future studies should also include controlled experiments where evaluators independently assess translation quality using both holistic judgment and formal scoring models, enabling direct comparison between model predictions and actual pass/fail outcomes.
8.2 Content-Type Specific Effects
Different content types (e.g., marketing, legal, technical, UI strings) may have different quality expectations and error tolerance. Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by client or use case.
We found that a single logarithmic formula performs well across a wide range of content types, suggesting that the logarithmic rule is quite universal—even with fixed parameters. However, this observation requires further validation through broader field data. A valuable extension of this work would be to investigate whether different genres (content types) exhibit distinct curve shapes—for example, steeper or flatter logarithmic profiles—or whether certain domains deviate significantly from the general trend. Such an analysis would help determine whether the proposed model serves as a general framework, as our experience suggests, or represents just one of several viable scoring paradigms.
We hypothesize a soft saturation at very large evaluation wordcount due to document‑level coherence expectations; future work can test whether $E(x)$ bends further (e.g., log‑log or logistic) at book/manual length.
8.3 Severity Weighting and Mixed Error Types
This paper concentrates on modeling tolerance in terms of penalty points, each of which typically represents one minor error. In practice, quality evaluations combine several error severities (minor, major, critical) penalized by specific multipliers and sometimes further scaled by different error-type weights applied for different content types.
Severity scaling (i.e., fixed severity weights such as minor/major/critical multipliers) is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points. This is separate from any error-type weights (e.g., terminology vs. grammar) a metric may also define. While a logarithmic curve fits minor-error tolerance well—minor errors being the most frequent and thus the most statistically stable—it may be less suitable for major errors, which occur far less often. For these rarer severities, an integer-valued sequence may be preferable, as indicated by Client 3’s data in Fig. 3. That sequence follows a loosely logarithmic trend but would require explicit rounding rules.
Future work should examine the form of such integer sequences, quantify how far they depart from a pure logarithmic curve, and determine the parameters that best describe severity scaling across different content types and languages.
8.4 Inter-Rater Reliability, IRR
When a linear model is used, the variability among human raters can be partly attributed to the fact that the model is only accurate at its anchoring point and loses precision elsewhere. In contrast, a calibrated non-linear model better reflects human judgment across a wide range of evaluation sample sizes. As a result, Inter-Rater Reliability (IRR) should increase significantly when using a non-linear calibrated scoring model, as it provides consistent and comparable measurements across different evaluation sample sizes, languages, and content types. More research is needed to quantify the improvement in inter-annotator agreement and to determine how much such a model reduces variance compared to traditional linear approaches in real-world annotation settings.
8.5 Extension to AI Evaluation Benchmarks
This paper also hypothesizes that the same principles of non-linear error perception apply to AI-generated text evaluation, including summarization, dialogue, and document translation by large language models. However, this claim has yet to be formally validated. AI benchmarks which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
Future work should involve testing the non-linear model as a normalization layer in LLM evaluation tasks, particularly those involving paragraph- or document-level generation, where holistic perception plays a major role.
A further implication is that a calibrated non-linear scoring model could eventually serve as a reference framework for future AI-based document-level evaluations, which currently provide holistic but qualitative judgments without a quantitative calibration.
9 Conclusion
Translation quality evaluation plays a critical role in ensuring that multilingual content meets the linguistic, technical, and functional standards required for global communication. While current industry practices predominantly rely on linear scoring models—extrapolating error thresholds proportionally from fixed reference wordcounts—this paper builds on empirical and practitioner evidence showing that linear models do not consistently align with human evaluation, motivating the adoption of non-linear alternatives.
This paper presents empirical data from experienced quality managers at large translation buyers, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size. This effect was evident across organizations, content types, and error categories and severities. This paper further supports these observations with theoretical underpinnings from psychophysics (via the Weber–Fechner law) and cognitive science (via Cognitive Load Theory), both of which explain why perceived quality and impact of errors do not scale linearly with text length.
To address this disconnect between the existing scoring systems and human judgment, this paper proposes a calibrated logarithmic scoring model that reflects the genuine growth of error tolerance. This model is mathematically simple, empirically grounded, and easily adaptable to specific client requirements. It provides a flexible, yet rigorous alternative to traditional linear error density calculations.
Because it retains a simple two-parameter form—and requires only a one-dimensional solve for calibration—the model is straightforward to integrate into CAT tools, dashboards, and analytics, while providing a unified way to reflect how perceived impact scales with error accumulation.
The implications of adopting this model are wide-ranging. Quality managers can gain a more accurate tool for defining passing thresholds in jobs of variable lengths. CAT tools and LQA systems and tools can offer context-aware scoring and fairer quality metrics. Evaluators benefit from consistency that reflects actual perception, and the same principles can be extended to evaluating AI-generated language output.
While further research is needed to validate and refine this model, particularly across languages, domains, content types, the fundamental insight seems to be clear: human perception of translation quality is inherently non-linear. By embracing this fact, the translation industry can take a critical step toward a more accurate, fair, and human-perception-aligned quality evaluation.
Finally, this paper reiterates the warning from [GladkoffEtAl2024]: below the sample size of approximately 250 words, no deterministic tolerance curve is statistically reliable. In such micro-range cases, binomial risk-based Statistical Quality Control (SQC) remains the only viable method. [^7] Even with SQC, micro-range evaluations cannot provide statistically reliable estimates of the underlying population quality, because the confidence intervals around error density are too wide to be meaningful at such small evaluation sample sizes. SQC explicitly departs from the idea of estimating population quality and instead evaluates the producer’s risk (the risk of wrongly rejecting acceptable output) and the consumer’s risk (the risk of wrongly accepting defective output). In this way, SQC reframes quality control for very small samples: It does not reduce statistical uncertainty, but it makes risks explicit and quantifiable for decision-making.
This study consolidates the multi-range proposal of [GladkoffEtAl2024] by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.
Acknowledgments
The authors gratefully acknowledge the MQM Council and the ASTM WK46396 Working Group (MQM 2.0: Analytic Translation Quality Evaluation) for providing a collaborative forum in which earlier versions of the non-linear scoring model were discussed, refined, and incorporated into the broader MQM framework. In particular, the authors would like to thank Ingemar Strandvik (MQM Council) and Angelika Vasaa, Head of the Quality Coordination Unit at the European Parliament, for their insightful comments, institutional perspective, and many valuable discussions that helped shape the ideas presented in this paper. Views and opinions expressed are however their own of the authors and acknowledged contributors, and do not necessarily reflect those of the European Union or Parliament. Neither the European Union nor the Parliament’s authority can be held responsible for them.
The authors thank the Quality Managers of the participating organizations (referred to in this document as Clients C1–C3) for their generous contributions of time, insight, and real-world data. Their candid responses and openness to critical reflection were essential to this research. Gratitude is also expressed to the broader translation quality evaluation community for ongoing discussions that continue to refine our shared understanding of what quality means in multilingual communication.
Appendix A. Calibration from Two Points (Numerical)
Checking the model feasibility.
Given two tolerance points $(x_{0},E_{0})$, $(x_{1},E_{1})$ with $x_{0}\neq x_{1}$, $E_{0},E_{1}>0$, define $r=E_{1}/E_{0}$, $\rho=x_{1}/x_{0}$. A unique calibration with $a>0,b>0$ exists iff
$\min\{1,\rho\}<r<\max\{1,\rho\}.$
One‑dimensional root‑finding. Let $f(b)=\ln(1+bx_{1})-r\ln(1+bx_{0})$. Near the origin, $f(b)\approx b(x_{1}-rx_{0})$. As $b\to\infty$, $f(b)\sim(1-r)\ln b+\ln x_{1}-r\ln x_{0}$, which flips sign relative to the origin when the feasibility condition holds.
Bisection (robust).
  1. 1.
    Compute $r=E_{1}/E_{0}$ and verify feasibility.
  2. 2.
    Set $b_{\mathrm{lo}}\leftarrow 0^{+}$ and evaluate $f(b_{\mathrm{lo}})$.
  3. 3.
    Grow $b_{\mathrm{hi}}$ geometrically until $f(b_{\mathrm{lo}})\,f(b_{\mathrm{hi}})<0$.
  4. 4.
    Bisect until convergence: $b\leftarrow(b_{\mathrm{lo}}+b_{\mathrm{hi}})/2$ and update the bracket by the sign of $f(b)$.
  5. 5.
    Output $b^{\star}$ and $a^{\star}=E_{0}/\ln(1+b^{\star}x_{0})$.
Notes. (i) Changing the log base rescales $a$ only; we use the natural logarithm $\ln$. (ii) For $b\approx 0$, $\ln(1+bx)\approx bx$, so the model reduces locally to a linear rule, clarifying why a single linear slope works only near its reference sample size. (iii) With $n\geq 3$ points $\{(x_{i},E_{i})\}$, fit $(a,b)$ by nonlinear least squares under $a>0,b>0$.
Appendix B. Least-Squares Calibration from Multiple Points
Problem.
Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ with $x_{i}>0$ and $E_{i}>0$, estimate $a>0,b>0$ in
$E(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr)$
by (constrained) least squares:
$\min_{a>0,\;b>0}\;S(a,b)\;\;:=\;\;\sum_{i=1}^{n}\Bigl[E_{i}-a\,L_{i}(b)\Bigr]^{2},\qquad L_{i}(b):=\ln\!\bigl(1+b\,x_{i}\bigr).$
Profiling out $a$.
For any fixed $b>0$, the minimizer in $a$ is
$a(b)\;=\;\frac{\sum_{i}E_{i}\,L_{i}(b)}{\sum_{i}L_{i}^{2}(b)}.$
Substituting yields the one‑dimensional profiled objective
$S(b)\;=\;\sum_{i}E_{i}^{2}\;-\;\frac{\Bigl(\sum_{i}E_{i}\,L_{i}(b)\Bigr)^{\!2}}{\sum_{i}L_{i}^{2}(b)}\;=\;S_{0}-\frac{S_{1}(b)^{2}}{S_{2}(b)},$
where $S_{0}=\sum_{i}E_{i}^{2}$, $S_{1}(b)=\sum_{i}E_{i}\,L_{i}(b)$, and $S_{2}(b)=\sum_{i}L_{i}^{2}(b)$.
Derivative for 1‑D solvers (optional).
Let $L_{i}^{\prime}(b)=\dfrac{x_{i}}{1+b\,x_{i}}$. Then
$S_{1}^{\prime}(b)=\sum_{i}E_{i}\,L_{i}^{\prime}(b),\qquad S_{2}^{\prime}(b)=2\sum_{i}L_{i}(b)\,L_{i}^{\prime}(b),$
and
$S^{\prime}(b)\;=\;-\,\frac{2\,S_{1}(b)\,S_{1}^{\prime}(b)\,S_{2}(b)\;-\;S_{1}(b)^{2}\,S_{2}^{\prime}(b)}{S_{2}(b)^{2}}.$
While not strictly required, $S^{\prime}(b)$ enables Newton/Brent methods. In practice, $S(b)$ is unimodal over $b>0$ for the elicited data we observed.
Numerically stable 1‑D procedure.
  1. 1.
    If a feasible two‑point calibration is available ( Appendix A. Calibration from Two Points (Numerical)), use its $b^{\star}$ as the starting value; otherwise set $b_{0}=1/(10\,\max_{i}x_{i})$.
  2. 2.
    Define $q(b)=S(b)$. Bracket a minimum by expanding geometrically from $b_{0}$ (e.g., multiply by $2$) until $q$ increases on both sides.
  3. 3.
    Minimize $q(b)$ on the bracket with a derivative‑free method (Brent or golden‑section).
  4. 4.
[…]
Weighted least squares (optional).
If different points have different reliabilities, use positive weights $w_{i}$ to minimize
$S_{w}(a,b)=\sum_{i}w_{i}\bigl[E_{i}-a\,L_{i}(b)\bigr]^{2}$
with
$a_{w}(b)=\frac{\sum_{i}w_{i}E_{i}\,L_{i}(b)}{\sum_{i}w_{i}L_{i}^{2}(b)},\qquad S_{w}(b)=\sum_{i}w_{i}E_{i}^{2}-\frac{\bigl(\sum_{i}w_{i}E_{i}L_{i}(b)\bigr)^{2}}{\sum_{i}w_{i}L_{i}^{2}(b)}.$
Then proceed as above with $S_{w}(b)$.
Approximate standard errors.
[…]
$\frac{\partial r_{i}}{\partial a}=-L_{i}(\hat{b}),\qquad\frac{\partial r_{i}}{\partial b}=-\,\hat{a}\,L_{i}^{\prime}(\hat{b})\;=\;-\,\hat{a}\,\frac{x_{i}}{1+\hat{b}\,x_{i}}.$
[…]
$\widehat{\mathrm{Cov}}(\hat{a},\hat{b})\;=\;\hat{\sigma}^{2}\,(J^{\top}J)^{-1}.$
For any $x>0$, the delta method gives the variance of predicted tolerance
$\widehat{\mathrm{Var}}\!\left[E(x)\right]\;\approx\;\nabla_{\theta}E(x)^{\top}\,\widehat{\mathrm{Cov}}(\hat{a},\hat{b})\,\nabla_{\theta}E(x),\quad\nabla_{\theta}E(x)=\begin{bmatrix}\ln(1+\hat{b}\,x)\\[3.0pt] \hat{a}\,\dfrac{x}{1+\hat{b}\,x}\end{bmatrix}$.
A nonparametric alternative is a bootstrap over the elicited points.
Implementation notes.
(i) Use $\ln$ for the natural log and enforce $b>0$. (ii) Rescale $x$ if needed to keep $1+\hat{b}\,x$ well‑conditioned numerically. (iii) When points are very near two‑point feasible, the LS minimum occurs close to the two‑point solution.
[…]
Sensitivity to anchors (illustrative). With the second point fixed at $(x_{1},E_{1})=(250,2)$, varying the primary anchor by $\pm 1$ at $x_{0}=1000$ yields:
Anchor $E_{0}$ $E(1000)$ $E(2000)$ $E(3000)$
4 4.00 5.16 5.86
5 (baseline) 5.00 7.05 8.36
6 6.00 9.30 11.59
Near the anchor, predictions are stable by construction; farther away, two‑point calibrations diverge—hence the recommendation to use least squares with multiple points when available (this Appendix B).
Appendix C. Excel Goal Seek method for Two-Point Calibration
  1. 1.
    Put $x_{0}$ in A1, $E_{0}$ in B1, $x_{1}$ in C1, and $E_{1}$ in D1.
  2. 2.
    Enter an initial guess for $b>0$ in E1.
  3. 3.
    Compute $\varphi(b)$ in cell F1 using the formula:
    =LN(1 + \$E\$1 * \$C\$1) - (\$D\$1 / \$B\$1) * LN(1 + \$E\$1 * \$A\$1)
  4. 4.
    Go to Data $\rightarrow$ What-If Analysis $\rightarrow$ Goal Seek: set cell F1 to 0 by changing cell E1.
  5. 5.
    Once Goal Seek converges, compute $a$ in cell G1 using the formula:
    =\$B\$1 / LN(1 + \$E\$1 * \$A\$1)
Appendix D. Minimal Python code for Two-Point Calibration
Below is a minimal two-point calibration function that computes the non-linear tolerance parameters $a$ (overall scale) and $b$ (curvature) as described in the paper.
import math
def two_point_calibrate(x0, E0, x1, E1):
    r = E1 / E0
    def f(b):  # score-equation for b
        return math.log(1.0 + b*x1) - r*math.log(1.0 + b*x0)
    # bracket a root
    blo = 1e-12
    f_lo = f(blo)
    bhi = 1e-6
    for _ in range(80):
        f_hi = f(bhi)
        if f_lo * f_hi < 0:
            break
        bhi *= 2.0
    else:
        raise ValueError("Could not bracket a root for b; check inputs.")
    # bisection
    for _ in range(120):
        bmid = 0.5 * (blo + bhi)
        f_mid = f(bmid)
        if abs(f_mid) < 1e-12:
            blo = bhi = bmid
            break
        if f_lo * f_mid < 0:
            bhi = bmid
        else:
            blo = bmid
            f_lo = f_mid
    b = 0.5 * (blo + bhi)
    a = E0 / math.log(1.0 + b*x0)
    return a, b
Command-line usage: Save as calibrate.py and append the wrapper below to call it from a terminal:
if __name__ == "__main__":
    import argparse, math
    p = argparse.ArgumentParser(
        description="Two-point calibration for E(x)=a*ln(1+b*x)")
    p.add_argument("--x0", type=float, required=True, help="reference size x0")
    p.add_argument("--E0", type=float, required=True, help="tolerance at x0")
    p.add_argument("--x1", type=float, required=True, help="second size x1")
    p.add_argument("--E1", type=float, required=True, help="tolerance at x1")
    p.add_argument("--x",  type=float, help="optional size to evaluate E(x)")
    args = p.parse_args()
    a, b = two_point_calibrate(args.x0, args.E0, args.x1, args.E1)
    print(f"a={a:.6f}, b={b:.8g}")
    if args.x is not None:
        E = a * math.log(1.0 + b*args.x)
        print(f"E({args.x:g})={E:.6f}")
Example 1 (reproduce §5.3).
python3 calibrate.py --x0 1000 --E0 5 --x1 250 --E1 2
Expected output (approx.): a=3.688000, b=0.00288
Example 2 (compute E(3,000)E(3{,}000), §7.3).
python3 calibrate.py --x0 1000 --E0 5 --x1 250 --E1 2 --x 3000
Expected output (approx.): E(3000)=8.357
Appendix E. Least-Squares Fit (multiple points)
This routine fits the logarithmic tolerance $E(x)=a\,\ln(1+bx)$ to multiple tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ by profiling out $a$ and minimizing the one-dimensional objective in $b$ (see App. B). Given any $b>0$,
$a(b)=\frac{\sum_{i}E_{i}\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}\!,\quad S(b)=\sum_{i}\!\left(E_{i}-a(b)\ln(1+bx_{i})\right)^{2},$
[…]
Python code for Least Squares Fit:
import math
def calibrate_log_lsq(points, weights=None, b0=None, grow=2.0, tol=1e-8):
    Least-squares fit of E(x) = a * ln(1 + b x) to multiple points.
    points  : list of (x, E) pairs with x>0, E>0
    weights : optional list of positive weights w_i (same length as points)
    b0      : optional initial guess for b; default 1 / (10 * max x)
    returns : (a_hat, b_hat, sse)
    xs, Es = zip(*points)
    n = len(xs)
    if weights is None:
        ws = [1.0] * n
    else:
        if len(weights) != n:
            raise ValueError("weights must match points length")
        ws = list(weights)
    xmax = max(xs)
    if b0 is None or b0 <= 0:
        b0 = 1.0 / (10.0 * xmax)
    def L(b):
        # vector of ln(1 + b x_i); return None if invalid
        try:
            return [math.log(1.0 + b * x) for x in xs]
        except ValueError:
            return None
    def a_of_b(b):
        Li = L(b)
        if Li is None or min(1.0 + b * x for x in xs) <= 0:
            return None
        num = sum(w * E * l for w, E, l in zip(ws, Es, Li))
        den = sum(w * l * l for w, l in zip(ws, Li))
        if den <= 0:
            return None
        return num / den
    def sse_profile(b):
        a = a_of_b(b)
        if a is None:
            return float("inf")
        Li = L(b)
        r = [math.sqrt(w) * (E - a * l) for w, E, l in zip(ws, Es, Li)]
        return sum(ri * ri for ri in r)
    # ---- bracket a minimum around b0 by geometric expansion
    a_l = b0 / grow
    a_c = b0
    a_r = b0 * grow
    f_l = sse_profile(a_l)
    f_c = sse_profile(a_c)
    f_r = sse_profile(a_r)
    # ensure f_c is the smallest
    it = 0
    while not (f_c <= f_l and f_c <= f_r) and it < 80:
        it += 1
        if f_l < f_r:
            a_r, f_r = a_c, f_c
            a_c, f_c = a_l, f_l
            a_l /= grow
            f_l = sse_profile(a_l)
        else:
            a_l, f_l = a_c, f_c
            a_c, f_c = a_r, f_r
            a_r *= grow
            f_r = sse_profile(a_r)
    if it == 80:
        raise ValueError("Could not bracket a minimum for b; check data.")
    # ---- golden-section search on [a_l, a_r]
    phi = (math.sqrt(5) - 1) / 2
    left, right = a_l, a_r
    x1 = right - phi * (right - left)
    x2 = left + phi * (right - left)
    f1 = sse_profile(x1)
    f2 = sse_profile(x2)
    while (right - left) > tol * (abs(left) + abs(right) + 1.0):
        if f1 > f2:
            left = x1
            x1, f1 = x2, f2
            x2 = left + phi * (right - left)
            f2 = sse_profile(x2)
        else:
            right = x2
            x2, f2 = x1, f1
            x1 = right - phi * (right - left)
            f1 = sse_profile(x1)
    b_hat = 0.5 * (left + right)
    a_hat = a_of_b(b_hat)
    sse = sse_profile(b_hat)
    return a_hat, b_hat, sse
To test this code from the command line, add the wrapper below (or place it in a separate lsq_fit.py) to fit from a list of $(x,E)$ pairs and optionally evaluate $E(x)$ at a target size.
if __name__ == "__main__":
    import argparse, math, sys
    p = argparse.ArgumentParser(
        description="Least-squares fit for E(x)=a*ln(1+b*x)")
    p.add_argument("--xy", action="append", metavar="X,E",
                   help="data point as ’x,E’ (repeat for multiple points)")
    p.add_argument("--x", type=float, help="optional size to evaluate E(x)")
    args = p.parse_args()
    if not args.xy:
        sys.exit("Provide at least one --xy ’x,E’ pair.")
    pts = []
    for pair in args.xy:
        x_str, E_str = pair.split(",")
        pts.append((float(x_str), float(E_str)))
    a, b, sse = calibrate_log_lsq(pts)
    print(f"a={a:.6f}, b={b:.8g}, SSE={sse:.6f}")
    if args.x is not None:
        print(f"E({args.x:g})={a*math.log(1.0+b*args.x):.6f}")
Example 1 (Client 1, pages vs. minor errors; Fig. 3/Table 1).
python3 lsq_fit.py \
  --xy 2,2 --xy 3,3 --xy 4,4 --xy 5,5 --xy 7,6 --xy 10,7 --xy 20,8
[…]
Example 2 (evaluate the fitted curve at 12 pages).
python3 lsq_fit.py \
  --xy 2,2 --xy 3,3 --xy 4,4 --xy 5,5 --xy 7,6 --xy 10,7 --xy 20,8 \
  --x 12
[…]
Appendix F. Wolfram Notebook for Least Squares Fit
This appendix provides a Wolfram Notebook implementation of the least-squares calibration method described in Section 5.3 and detailed in Appendix A. The code can be pasted directly into a Wolfram Notebook and run cell by cell. It reproduces the profiled least-squares fitting procedure for the model
$E(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr),\qquad a>0,\;b>0,$
using bracketing and golden-section search over $b$, with $a$ profiled in closed form.
Instructions
  1. 1.
    Open Wolfram Mathematica and create a new Notebook.
  2. 2.
    Copy the code blocks below into separate cells.
  3. 3.
    Evaluate the cells in order (Shift+Enter).
  4. 4.
    The final cell produces a plot and fitted formula.
Cell 1: Definitions
ClearAll[CalibrateLogLSQ];
Options[CalibrateLogLSQ] = {"Weights"->None,"InitialB"->Automatic,
   "BracketStep"->2.,"Tolerance"->1.*^-8};
CalibrateLogLSQ[data_List, opts:OptionsPattern[]] := Module[
  {x,E,n,wOpt,w,b0,step,tol,L,Lp,aOfB,sseProfile,bracketMin,
   goldenMin,br,bHat,aHat,yHat,resid,sse,sigma2,J,JTJ,cov},
  x=N[data[[All,1]]]; E=N[data[[All,2]]]; n=Length[x];
  If[AnyTrue[x,#<=0&]||AnyTrue[E,#<=0&],Return[\$Failed,Module]];
  wOpt=OptionValue["Weights"];
  w=If[wOpt===None,ConstantArray[1.,n],N[wOpt]];
  If[Length[w]=!=n||AnyTrue[w,#<=0&],Return[\$Failed,Module]];
  b0=Replace[OptionValue["InitialB"],Automatic->(1./(10. Max[x]))];
  step=OptionValue["BracketStep"]; tol=OptionValue["Tolerance"];
  L[b_?NumericQ]:=Log[1.+b*x];
  Lp[b_?NumericQ]:=x/(1.+b*x);
  aOfB[b_?NumericQ]:=Module[{Li=L[b],num,den},
    If[Min[1.+b*x]<=0,Return[Indeterminate]];
    num=Total[w*E*Li]; den=Total[w*Li*Li];
    If[den<=0||!NumericQ[den],Indeterminate,num/den]];
  sseProfile[b_?NumericQ]:=Module[{a=aOfB[b],Li,r},
    If[!NumericQ[a],Return[Infinity]];
    Li=L[b]; r=Sqrt[w]*(E-a*Li); r.r];
  bracketMin[f_,bStart_,s_:2.,maxIter_:60]:=Module[{a,b,c,fa,fb,fc,it=0},
    a=bStart/s; b=bStart; c=bStart*s; fa=f[a]; fb=f[b]; fc=f[c];
    While[!(fb<fa&&fb<fc)&&it<maxIter,
      it++;
      If[fa<fb,
        c=b;fc=fb; b=a;fb=fa; a=a/s; If[a<=0,a=b/(2 s)]; fa=f[a],
        a=b;fa=fb; b=c;fb=fc; c=c*s; fc=f[c]]];
    {a,b,c}];
  goldenMin[f_,a0_,c0_,t_:1.*^-8,maxIter_:200]:=Module[
    {phi=(Sqrt[5]-1)/2.,a=a0,c=c0,x1,x2,f1,f2,it=0},
    x1=c-phi(c-a); x2=a+phi(c-a); f1=f[x1]; f2=f[x2];
    While[(c-a)>t(Abs[a]+Abs[c]+1.)&&it<maxIter,
      it++;
      If[f1>f2, a=x1; x1=x2; f1=f2; x2=a+phi(c-a); f2=f[x2],
                 c=x2; x2=x1; f2=f1; x1=c-phi(c-a); f1=f[x1]]];
    .5(a+c)];
  br=bracketMin[sseProfile,b0,step];
  bHat=goldenMin[sseProfile,br[[1]],br[[3]],tol];
  aHat=aOfB[bHat]; If[!NumericQ[aHat],Return[\$Failed,Module]];
  yHat=aHat*L[bHat]; resid=E-yHat;
  sse=(Sqrt[w]*resid).(Sqrt[w]*resid);
  sigma2=sse/Max[n-2,1];
  J=Transpose[{-L[bHat],-aHat*Lp[bHat]}];
  JTJ=If[wOpt===None,Transpose[J].J,Transpose[Sqrt[w]*J].(Sqrt[w]*J)];
  cov=sigma2*PseudoInverse[JTJ];
  <|"a"->aHat,"b"->bHat,"SSE"->sse,"Sigma2"->sigma2,
    "Residuals"->resid,"Covariance"->cov,
    "Predict"->(aHat*Log[1.+bHat*#]&),"Bracket"->br[[{1,3}]]|>
Cell 2: Fit and Quick Results
data = {{2,2},{3,3},{4,4},{5,5},{7,6},{10,7},{20,8}};
fit  = CalibrateLogLSQ[data];
{fit["a"], fit["b"], fit["SSE"]} // N
Cell 3: Plot
 With[{a = fit["a"], b = fit["b"]},
 Show[
     ListPlot[data, PlotStyle -> {Blue, PointSize[.02]}],
     ListLinePlot[data, InterpolationOrder -> 1,
                  PlotStyle -> {Blue, Dotted, Thick}],
     Plot[a*Log[1 + b*x], {x, 0.5, 25}, PlotStyle -> {Red, Thick}]
   GridLines -> Automatic, Frame -> True,
   FrameLabel -> {"Pages (x)", "Tolerance  E(x)"},
   PlotLabel -> Style["Least-squares log fit", 14, Bold],
  Epilog -> {
  Inset[
    Style[
      Row[{"E(x) \[TildeTilde] ", NumberForm[a, {5,5}], " ln(1 + ",
           NumberForm[b, {5,5}], " x)"}],
      12, Bold, Red],
    Scaled[{0.6, 0.65}]
   ImageSize -> Large
Illustration and Resulting Formula
Figure 8 shows the illustration produced by Cell 3. The fitted formula is returned by {fit["a"], fit["b"]} in Cell 2, which evaluates numerically to
$E(x)\;\approx\;\hat{a}\,\ln\!\bigl(1+\hat{b}\,x\bigr),$
with coefficients $\hat{a},\hat{b}$ depending on the dataset.
For the sample dataset serving as example, the approximated function is:
$E(x)\;\approx\;3.35301\,\ln\!\bigl(1+0.59046\,x\bigr),$
with residual sum of squares (SSE) $\approx 1.55087$.
For different dataset change data array in Cell 2.
Figure 8: Least-squares logarithmic fit in Wolfram Notebook.